Solve the following system of equations using any algebraic method-
2x 3y=17
5x 6y=32

Found 2 solutions by macston, raphaelfundani:
Answer by macston(5194)

2x+3y=17
5x+6y=32 Multiply first equation by 2 and subtract
4x+6y=34
x=-2 ANSWER 1: x=-2
Back solve in one equation for y
2x+3y=17
2(-2)+3y=17
-4+3y=17 Add 4 to each side.
3y=21 Divide each side by 3
y=7 ANSWER 2: y=7
CHECK:
5x+6y=32
5(-2)+6(7)=32
-10+42=32
32=32

Answer by raphaelfundani(3)

Solve 2x 3y=17 and 5x 6y=32
lets have the same value of x
5(2x 3y=17)
2(5x 6y=32)
:- 10x 15y=35
-(10x 12y=64)
:- 3y=-29
y=-29/3
then chose one equesion and substitute for y
2x 3(-29/3)=17
2x-29=17
2x=17 29=46
2x=46
x=23
:- the value of x is 23 and y is -29/3.